Predict the world’s population size in 2040.

Year

1927

1960

1974

1987

1999

2012

2018

Population

2

3

4

5

6

7

7.6
Answer #1

1927 1960 1974 1987 1999 2012 2018 X.Y 3854 5880 7896 9935 11994 14084 15336.8 2 3713329 3841600 3896676 3948169 3996001 4048

y = -121.457 +0.064x

at x = 2040

y = -121.457 +0.064 * 2040

y= -121.457 +130.56

y= 9.103